# Optimistic Task Inference for Behavior Foundation Models

### v1

@@ Abstract
Behavior Foundation Models (BFMs) are capable of retrieving high-performing policy for any reward function specified directly at test-time, commonly referred to as zero-shot reinforcement learning (RL). While this is a very efficient process in terms of compute, it can be less so in terms of data: as a standard assumption, BFMs require computing rewards over a non-negligible inference dataset, assuming either access to a functional form of rewards, or significant labeling efforts. To alleviate these limitations, we tackle the problem of task inference purely through interaction with the environment at test-time. We propose OpTI-BFM, an optimistic decision criterion that directly models uncertainty over reward functions and guides BFMs in data collection for task inference. Formally, we provide a regret bound for well-trained BFMs through a direct connection to upper-confidence algorithms for linear bandits. Empirically, we evaluate OpTI-BFM on established zero-shot benchmarks, and observe that it enables successor-features-based BFMs to identify and optimize an unseen reward function in a handful of episodes with minimal compute overhead. Code is available at https://github.com/ThomasRupf/opti-bfm.
@@ 1 Introduction
Zero-shot reinforcement learning (Touati et al., 2023) has gradually gained relevance as a powerful generalization of standard, single-reward RL (Sutton & Barto, 2018; Silver et al., 2021). Zero-shot agents are designed to distill optimal policies for a set of reward functions from unlabeled, offline data (Touati & Ollivier, 2021; Agarwal et al., 2024; Park et al., 2024; Jajoo et al., 2025). As these methods scale to more complex and broader environments, they are often referred to as Behavior Foundation Models (BFMs) (Park et al., 2024; Tirinzoni et al., 2025). At their core, the majority of BFMs rely on Universal Successor Features (Ma et al. (2020), USFs). These methods build upon explicit representations of states (features): given a set of policies, the expected discounted sum of features along each policy’s trajectory may be estimated completely offline (e.g., through TD learning) (Dayan, 1993; Barreto et al., 2017). As long as a reward function lies within the span of features, zero-shot policy evaluation can be performed through a simple scalar product, which in turn enables zero-shot policy improvement, i.e., each policy is implicitly paired with a reward function and updated towards optimality (Touati & Ollivier, 2021).
Once this unsupervised pre-training phase is complete, USFs yield a set of learned policies. Given a reward function that is linear in the features, its optimal policy is by construction indexed by the linear weights describing the reward function in the basis of features, which can in turn be interpreted as task embeddings. The process of finding the optimal policy, which we refer to as task inference, thus corresponds to solving a linear regression problem. While this process is remarkably efficient in terms of compute, it retains strict requirements in terms of data: a dataset of labeled (state, reward) pairs needs to be provided. In simple settings, this necessity is not particularly problematic: BFMs have been largely trained in simulations (Touati et al., 2023; Park et al., 2024; Tirinzoni et al., 2025), which makes generating reward labels for the pre-training dataset, or additional data, particularly convenient. Realistically, however, (i) the pre-training dataset might be unavailable or proprietary and, most importantly, (ii) labeling states with rewards might incur significant costs. For instance, when BFMs are pre-trained directly from pixels, evaluating success from a single image is not a straightforward or cheap operation.
@@ 2 Preliminaries
A similar structure also holds for value functions: $V_{r}^{\pi}(s)=z^{\top}\psi^{\pi}(s)$, when defining $\psi^{\pi}(s)=\mathbb{E}_{a\sim\pi(\cdot|s)}\psi^{\pi}(s,a)$. Behavior Foundation Models (BFMs) generally [^1] There are BFMs that are not based on USFs. We refer the reader to Agarwal et al. (2025a) for a comprehensive overview. capitalize on the opportunity of evaluating policies for multiple reward functions, by additionally learning a family of parameterized policies $(\pi_{z})_{z\in\mathcal{Z}}$ with respect to all rewards in the span of features $\phi$ (Borsa et al., 2018; Touati & Ollivier, 2021; Park et al., 2024; Agarwal et al., 2025b). Concretely, BFMs train a family of parameterized policies so that each policy $\pi_{z}$ is optimal for the reward function $r(s)=\phi(s)^{\top}z$:
@@ 3.2 Method: OpTI-BFM
Update estimator $(\hat{z}_{n+t-1},V_{n+t-1})$ with $(\phi(s_{t}),r_{t})$ $\triangleright$ Update based on reward-feedback.
@@ 5.1 How does OpTI-BFM compare to other online task inference methods?
We first evaluate all methods in the online task inference setting described in Section˜3.1, and report episodic returns (Eq.˜5) in Fig.˜2. We find that OpTI-BFM recovers Oracle performance on all tasks within 5 episodes (5k environment steps) of interaction. We observe a significant gap between the optimistic strategy of OpTI-BFM and its TS variant in Cheetah. Nevertheless, TS remains a promising approach, as it avoids any optimization problem, and simply samples task embeddings from its current belief. Finally, we find that LoLA, which ignores the linear structure of the problem and performs blackbox policy search, makes slower progress which is better visible over 50 episodes in Fig.˜10 in Appendix˜C. This result is consistent with existing ablations initializing LoLA to a random task embedding (see Sikchi et al. (2025), Figure 5), as is the case in our setting.
Figure 2: Mean relative performance over 10 episodes of interaction in DMC. OpTI-BFM recovers Oracle performance in 5 episodes. We report per-task absolute performance in Fig.˜8 in Appendix˜C.
@@ 5.2 Is the data collected actively by OpTI-BFM informative?
Figure 3: Relative performance after different # of environment interactions. OpTI-BFM is consistently among the top performers for all environments and time-steps. We show per task performance in Fig.˜9 in Appendix˜C.
@@ 5.3 Are frequent updates important?
The regret bound provided in Section˜3.3 assumes a version of OpTI-BFM that only updates the task embedding it executes at the start of each episode. In this section we aim to experimentally quantify this gap in theory and practice by evaluating this episodic version of OpTI-BFM. Fig.˜4 shows that changing the latent on an episode-level leads to slower improvement than changing every step. The episodic versions do reach equal performance eventually as shown in Fig.˜10 in Appendix˜C.
Figure 4: Relative performance of our methods, and their variants that keep task embeddings fixed for each episode (-EP). Updating the task embedding during the episode leads to faster convergence. For longer episode evaluations see Fig.˜10 in Appendix˜C.
@@ 5.5 Can OpTI-BFM warm-start from labeled data?
Figure 6: Return of OpTI-BFM in DMC Cheetah when warm-starting with $n$ i.i.d. labeled states from the training dataset. Initial performance increases quickly as $n$ grows. We report full results in Fig.˜11 in Appendix˜C.
@@ 6 Related Work
This work builds upon USFs (Dayan, 1993; Barreto et al., 2017; Ma et al., 2020), and specifically on BFMs that utilize SFs to train a parameterized policy and the corresponding parameterized SFs. A common approach for learning SF-based BFMs is to approximate of successor measures (Blier et al., 2021; Touati & Ollivier, 2021; Agarwal et al., 2025b), but other options spanning from spectral decomposition of random visitations (Wu et al., 2018) to implicit value learning (Park et al., 2024) also exist. In this regard, we refer the reader to Agarwal et al. (2025a) for a broader overview of BFMs and unsupervised RL. Among these approaches, the Forward Backward (FB) framework (Touati & Ollivier, 2021) has risen to prominence, as it achieves state-of-the-art performance in standard continuous control tasks (Touati et al., 2023), while being scalable to humanoid control (Tirinzoni et al., 2025). Recent work has explored various applications of FB’s zero-shot capabilities, such as imitation learning (Pirotta et al., 2023), epistemic exploration (Urpí et al., 2025), and constrained RL (Hugessen et al., 2025). Among related works, our method is most closely related to Sikchi et al. (2025), which however focuses on a different problem: fine-tuning a zero-shot policy, which is already initialized through standard offline task inference. On the other hand, our goal is to find a strong policy purely through environment interaction, without relying on preexisting labeled data. The two approaches might nevertheless be combined.
@@ 7 Conclusion
OpTI-BFM is designed to perform task inference in a minimal number of episodes: this is mainly possible as the search space is minimal. While updating the task embedding alone enables great sample efficiency, fine-tuning additional components of the BFM may provide even better performance in the long run (Sikchi et al., 2025).
@@ A.4 Regret Bound
A.5 Useful Properties
A.5.1 Determinants

### v2

@@ Abstract
Behavior Foundation Models (BFMs) are capable of retrieving high-performing policy for any reward function specified directly at test-time, commonly referred to as zero-shot reinforcement learning (RL). While this is a very efficient process in terms of compute, it can be less so in terms of data: as a standard assumption, BFMs require computing rewards over a non-negligible inference dataset, assuming either access to a functional form of rewards, or significant labeling efforts. To alleviate these limitations, we tackle the problem of task inference purely through interaction with the environment at test-time. We propose OpTI-BFM, an optimistic decision criterion that directly models uncertainty over reward functions and guides BFMs in data collection for task inference. Formally, we provide a regret bound for well-trained BFMs through a direct connection to upper-confidence algorithms for linear bandits. Empirically, we evaluate OpTI-BFM on established zero-shot benchmarks, and observe that it enables successor-features-based BFMs to identify and optimize an unseen reward function in a handful of episodes with minimal compute overhead. [^1] Code is available at https://github.com/ThomasRupf/opti-bfm.
@@ 1 Introduction
Zero-shot reinforcement learning (Touati et al., 2023) has gradually gained relevance as a powerful generalization of standard, single-reward RL (Sutton & Barto, 2018; Silver et al., 2021). Zero-shot agents are designed to distill optimal policies for a set of reward functions from unlabeled, offline data (Touati & Ollivier, 2021; Agarwal et al., 2024; Park et al., 2024b; Jajoo et al., 2025). As these methods scale to more complex and broader environments, they are often referred to as Behavior Foundation Models (BFMs) (Park et al., 2024b; Tirinzoni et al., 2025). At their core, the majority of BFMs rely on Universal Successor Features (Ma et al. (2020), USFs). These methods build upon explicit representations of states (features): given a set of policies, the expected discounted sum of features along each policy’s trajectory may be estimated completely offline (e.g., through TD learning) (Dayan, 1993; Barreto et al., 2017). As long as a reward function lies within the span of features, zero-shot policy evaluation can be performed through a simple scalar product, which in turn enables zero-shot policy improvement, i.e., each policy is implicitly paired with a reward function and updated towards optimality (Touati & Ollivier, 2021).
Once this unsupervised pre-training phase is complete, USFs yield a set of learned policies. Given a reward function that is linear in the features, its optimal policy is by construction indexed by the linear weights describing the reward function in the basis of features, which can in turn be interpreted as task embeddings. The process of finding the optimal policy, which we refer to as task inference, thus corresponds to solving a linear regression problem. While this process is remarkably efficient in terms of compute, it retains strict requirements in terms of data: a dataset of labeled (state, reward) pairs needs to be provided. In simple settings, this necessity is not particularly problematic: BFMs have been largely trained in simulations (Touati et al., 2023; Park et al., 2024b; Tirinzoni et al., 2025), which makes generating reward labels for the pre-training dataset, or additional data, particularly convenient. Realistically, however, (i) the pre-training dataset might be unavailable or proprietary and, most importantly, (ii) labeling states with rewards might incur significant costs. For instance, when BFMs are pre-trained directly from pixels, evaluating success from a single image is not a straightforward or cheap operation.
@@ 2 Preliminaries
A similar structure also holds for value functions: $V_{r}^{\pi}(s)=z^{\top}\psi^{\pi}(s)$, when defining $\psi^{\pi}(s)=\mathbb{E}_{a\sim\pi(\cdot|s)}\psi^{\pi}(s,a)$. Behavior Foundation Models (BFMs) generally [^2] There are BFMs that are not based on USFs. We refer the reader to Agarwal et al. (2025a) for a comprehensive overview. capitalize on the opportunity of evaluating policies for multiple reward functions, by additionally learning a family of parameterized policies $(\pi_{z})_{z\in\mathcal{Z}}$ with respect to all rewards in the span of features $\phi$ (Borsa et al., 2018; Touati & Ollivier, 2021; Park et al., 2024b; Agarwal et al., 2025b). Concretely, BFMs train a family of parameterized policies so that each policy $\pi_{z}$ is optimal for the reward function $r(s)=\phi(s)^{\top}z$:
@@ 3.2 Method: OpTI-BFM
Update $(\hat{z}_{n+t-1},V_{n+t-1})$ with $(\phi(s_{t}),r_{t})$ through Eq. 8 $\triangleright$ Update based on reward-feedback.
@@ 5.1 How does OpTI-BFM compare to other online task inference methods?
We first evaluate all methods in the online task inference setting described in Section˜3.1, and report episodic returns (Eq.˜5) in Fig.˜2. We find that OpTI-BFM recovers Oracle performance on all tasks within 5 episodes (5k environment steps) of interaction. We observe a significant gap between the optimistic strategy of OpTI-BFM and its TS variant in Cheetah. Nevertheless, TS remains a promising approach, as it avoids any optimization problem, and simply samples task embeddings from its current belief. Finally, we find that LoLA, which ignores the linear structure of the problem and performs blackbox policy search, makes slower progress which is better visible over 50 episodes in Fig.˜16 in Appendix˜C. This result is consistent with existing ablations initializing LoLA to a random task embedding (see Sikchi et al. (2025), Figure 5), as is the case in our setting.
Figure 2: Mean relative performance over 10 episodes of interaction in DMC. OpTI-BFM recovers Oracle performance in 5 episodes. We report per-task absolute performance in Fig.˜14 in Appendix˜C.
@@ 5.2 Is the data collected actively by OpTI-BFM informative?
Figure 3: Relative performance after different # of environment interactions. OpTI-BFM is consistently among the top performers for all environments and time-steps. We show per task performance in Fig.˜15 in Appendix˜C.
@@ 5.3 Are frequent updates important?
The regret bound provided in Section˜3.3 assumes a version of OpTI-BFM that only updates the task embedding it executes at the start of each episode. In this section we aim to experimentally quantify this gap in theory and practice by evaluating this episodic version of OpTI-BFM. Fig.˜4 shows that changing the latent on an episode-level leads to slower improvement than changing every step. The episodic versions do reach equal performance eventually as shown in Fig.˜16 in Appendix˜C. Intuitively, this faster improvement may be explained by the fact that the default version of OpTI-BFM can adjust faster to new information.
Figure 4: Relative performance of our methods, and their variants that keep task embeddings fixed for each episode (-EP). Updating the task embedding during the episode leads to faster convergence. For longer episode evaluations see Fig.˜16 in Appendix˜C.
@@ 5.5 Can OpTI-BFM warm-start from labeled data?
Figure 6: Return of OpTI-BFM in DMC Cheetah when warm-starting with $n$ i.i.d. labeled states from the training dataset. Initial performance increases quickly as $n$ grows. We report full results in Fig.˜17 in Appendix˜C.
@@ 6 Related Work
This work builds upon USFs (Dayan, 1993; Barreto et al., 2017; Ma et al., 2020), and specifically on BFMs that utilize SFs to train a parameterized policy and the corresponding parameterized SFs. A common approach for learning SF-based BFMs is to approximate of successor measures (Blier et al., 2021; Touati & Ollivier, 2021; Agarwal et al., 2025b), but other options spanning from spectral decomposition of random visitations (Wu et al., 2018) to implicit value learning (Park et al., 2024b) also exist. In this regard, we refer the reader to Agarwal et al. (2025a) for a broader overview of BFMs and unsupervised RL. Among these approaches, the Forward Backward (FB) framework (Touati & Ollivier, 2021) has risen to prominence, as it achieves state-of-the-art performance in standard continuous control tasks (Touati et al., 2023), while being scalable to humanoid control (Tirinzoni et al., 2025). Recent work has explored various applications of FB’s zero-shot capabilities, such as imitation learning (Pirotta et al., 2023), epistemic exploration (Urpí et al., 2025), and constrained RL (Hugessen et al., 2025). Among related works, our method is most closely related to Sikchi et al. (2025), which however focuses on a different problem: fine-tuning a zero-shot policy, which is already initialized through standard offline task inference. On the other hand, our goal is to find a strong policy purely through environment interaction, without relying on preexisting labeled data. The two approaches might nevertheless be combined.
@@ 7 Conclusion
OpTI-BFM is designed to perform task inference in a minimal number of episodes: this is mainly possible as the search space is minimal. While updating the task embedding alone enables great sample efficiency, fine-tuning additional components of the BFM may provide even better performance in the long run (Sikchi et al., 2025). Moreover, our theoretical guarantees only cover slower, episode-level updates: extending these results to (empirically stronger) per-step represents an important direction for future formal works.
@@ A.4 Regret Bound
A.5 Regret Bound under Weaker Assumptions
Our empirical evaluations are performed in a setting which violates two assumptions described in A.1.1: perfect USF estimation (A1) and linearity of rewards (A2). This section provides a study of the algorithm’s behavior under these violations taking some ideas from existing theory of misspecified bandits (Ghosh et al., 2017; Bogunovic & Krause, 2021; Lattimore & Szepesvári, 2020).
We start by replacing A1 and A2 by weaker assumptions that quantify the mismatch between $\phi$ and $\psi$ and between $r$ and its projection $z_{r}^{\top}\psi$ respectively:
  • (A1’)
    For every $(s,a,z)\in\mathcal{S\times A\times Z}$ we have
    $\left\lVert{\psi(s,a,z)-\textstyle{\sum}_{t=0}^{H-1}\gamma^{t}\mathbb{E}\left[{\phi(s_{t})}\ \middle|\ {\pi_{z},s,a}\right]}\right\rVert_{2}\leq\xi$ (71)
    and
  • (A2’)
    For reward $r$ we have for each state $s_{k,t}$
    $\mathbb{E}[\|r(s_{k,t})-\phi(s_{k,t})^{\top}z_{r}\|]\leq\zeta$ (73)
    where the expectation is over i.i.d. mean-zero $\sigma$-subgaussian noise $\eta_{k,t}$.
We may thus write $\psi$ as a biased version of the true $\psi^{\star}$
$\psi(s,a,z)+\Delta_{\psi}(s,a,z)=\psi^{*}(s,a,z),$ (74)
and, similarly, rewrite rewards as
$r_{k,t}=\phi(s_{k,t})^{\top}z_{r}+\Delta_{r}(s_{k,t})+\eta_{k,t}.$ (75)
We can then establish a uniform-in-time bound as in Theorem˜1 by inflating $\beta$ to account for the newly introduced bias in the reward. Considering the distance between the current mean estimate $\hat{z}_{t-1}$ and $z_{r}$, we can decompose this into
$\left\lVert{z_{r}-\hat{z}_{t-1}}\right\rVert_{V_{t-1}}$ $=\left\lVert{z_{r}-\sum_{i=1}^{t-1}\phi_{i}\phi_{i}^{\top}z_{r}-V_{t-1}^{-1}\sum_{i=1}^{t-1}\phi_{i}\eta_{i}-V_{t-1}^{-1}\sum_{i=1}^{t-1}\phi_{i}\Delta_{r}(s_{i})}\right\rVert_{V_{t-1}}$ (76)
$=\left\lVert{z_{r}-V_{t-1}z_{r}+\lambda z_{r}-V_{t-1}^{-1}\sum_{i=1}^{t-1}\phi_{i}\eta_{i}-V_{t-1}^{-1}\sum_{i=1}^{t-1}\phi_{i}\Delta_{r}(s_{i})}\right\rVert_{V_{t-1}}$ (77)
$=\left\lVert{z_{r}-z_{r}-V_{t-1}^{-1}\left(\lambda z_{r}+\sum_{i=1}^{t-1}\phi_{i}\eta_{i}+\sum_{i=1}^{t-1}\phi_{i}\Delta_{r}(s_{i})\right)}\right\rVert_{V_{t-1}}$ (78)
$=\left\lVert{\lambda z_{r}+\sum_{i=1}^{t-1}\phi_{i}\eta_{i}+\sum_{i=1}^{t-1}\phi_{i}\Delta_{r}(s_{i})}\right\rVert_{V_{t-1}^{-1}}$ (79)
$\leq\left\lVert{\lambda z_{r}+\sum_{i=1}^{t-1}\phi_{i}\eta_{i}}\right\rVert_{V_{t-1}^{-1}}+\left\lVert{\sum_{i=1}^{t-1}\phi_{i}\Delta_{r}(s_{i})}\right\rVert_{V_{t-1}^{-1}}$ (80)
The first term is covered by Theorem˜1, while the new bias term can be bounded through similar techniques to those used in the main proof:
$\left\lVert{\sum_{i=1}^{t-1}\phi_{i}\Delta_{r}(s_{i})}\right\rVert_{V_{t-1}^{-1}}$ $\leq\sum_{i=1}^{t-1}|\Delta_{r}(s_{i})|\left\lVert{\phi_{i}}\right\rVert_{V_{t-1}^{-1}}$ $\left\lVert{\cdot}\right\rVert_{V_{t-1}^{-1}}\text{ properties}$ (81)
$\leq\zeta\sum_{i=1}^{t-1}\left\lVert{\phi_{i}}\right\rVert_{V_{t-1}^{-1}}$ A2’ (82)
$\leq\zeta\sum_{i=1}^{t-1}\left\lVert{\phi_{i}}\right\rVert_{V_{i-1}^{-1}}$ $V_{i-1}\preccurlyeq V_{t-1}$ (83)
$\leq\zeta\sqrt{t\sum_{i=1}^{t-1}\left\lVert{\phi_{i}}\right\rVert^{2}_{V_{i-1}^{-1}}}$ Cauchy-Schwarz (84)
$\leq\zeta\sqrt{t\sum_{i=1}^{t-1}\min\{\tfrac{L^{2}}{\lambda},\left\lVert{\phi_{i}}\right\rVert^{2}_{V_{i-1}^{-1}}\}}$ (85)
$\leq\frac{L\zeta}{\sqrt{\lambda}}\sqrt{t\sum_{i=1}^{t-1}\min\{1,\left\lVert{\phi_{i}}\right\rVert^{2}_{V_{i-1}^{-1}}\}}$ (86)
$\leq\frac{\sqrt{2}L\zeta}{\sqrt{\lambda}}\sqrt{t\log\left(\frac{\det V_{t}}{\det V_{0}}\right)}$ (87)
As a result, we can establish a uniform-in-time guarantee if $\beta$ is inflated as
$\beta_{t}^{\prime}=\beta_{t}+\frac{\sqrt{2}L\zeta}{\sqrt{\lambda}}\sqrt{t\log\left(\frac{\det V_{t}}{\det V_{0}}\right)}.$ (88)
Continuing to build the regret bound, we first notice that the mismatch also impacts the discounted return:
$G_{k,z}$ $=\mathbb{E}\left[{\hat{G}_{k}}\ \middle|\ {\pi_{z},s_{k,0}}\right]$ (89)
$\stackrel{{\scriptstyle A2^{\prime}}}{{=}}\mathbb{E}\left[{\textstyle{\sum}_{t=0}^{H-1}\gamma^{t}\langle z_{r},\phi(s_{k,t})\rangle+\Delta_{r}(s_{k,t})}\ \middle|\ {\pi_{z},s_{k,0}}\right]$ (91)
$=\left\langle z_{r},\mathbb{E}\left[{\textstyle{\sum}_{t=0}^{H-1}\gamma^{t}\phi(s_{k,t})}\ \middle|\ {\pi_{z},s_{k,0}}\right]\right\rangle+\mathbb{E}\left[{\textstyle{\sum}_{t=0}^{H-1}\gamma^{t}\Delta_{r}(s_{k,t})}\ \middle|\ {\pi_{z},s_{k,0}}\right]$ (92)
$=\langle z_{r},\psi^{*}(s_{k,0},z)\rangle+\mathbb{E}\left[{\textstyle{\sum}_{t=0}^{H-1}\gamma^{t}\Delta_{r}(s_{k,t})}\ \middle|\ {\pi_{z},s_{k,0}}\right]$ (93)
$\stackrel{{\scriptstyle A2^{\prime}}}{{\leq}}\langle z_{r},\psi^{*}(s_{k,0},z)\rangle+\tfrac{\zeta}{1-\gamma}$ (94)
We can then proceed along the main proof in Appendix A.4 as normal, substituting $\beta^{\prime}$ for $\beta$ throughout. Given the suboptimal $\psi$ Eq.˜43 now becomes
$\langle\psi^{\star}(s_{k,0},z_{r}),z_{r}\rangle\leq\text{UCB}_{k}(z_{r})\leq\text{UCB}_{k}(z_{k})=\langle\psi^{\star}(s_{k,0},z_{k}),\tilde{z}_{k}\rangle.$ (95)
Then, in the bound on the instantaneous regret (Eq.˜44), we get an added term.
$e_{k}$ $=G_{k,z_{r}}-G_{k,z_{k}}$ (96)
$\leq\langle\psi^{\star}(s_{k,0},z_{r}),z_{r}\rangle-\langle\psi^{\star}(s_{k,0},z_{k}),z_{r}\rangle+\tfrac{2\zeta}{1-\gamma}$ (97)
$\leq\langle\psi^{\star}(s_{k,0},z_{k}),\tilde{z}_{k}\rangle-\langle\psi^{\star}(s_{k,0},z_{k}),z_{r}\rangle+\tfrac{2\zeta}{1-\gamma}$ (98)
$=\langle\psi^{\star}(s_{k,0},z_{k}),\tilde{z}_{k}-z_{r}\rangle+\tfrac{2\zeta}{1-\gamma}$ (99)
$\stackrel{{\scriptstyle A2^{\prime}}}{{=}}\langle\psi(s_{k,0},z_{k}),\tilde{z}_{k}-z_{r}\rangle+\langle\Delta_{\psi}(s_{k,0},z_{k}),\tilde{z}_{k}-z_{r}\rangle+\tfrac{2\zeta}{1-\gamma}$ (100)
$\leq\langle\psi(s_{k,0},z_{k}),\tilde{z}_{k}-z_{r}\rangle+\left\lVert{\Delta_{\psi}(s_{k,0},z_{k})}\right\rVert_{V_{k-1}^{-1}}\left\lVert{\tilde{z}_{k}-z_{r}}\right\rVert_{V_{k-1}}+\tfrac{2\zeta}{1-\gamma}$ Cauchy-Schwarz
$\leq\langle\psi(s_{k,0},z_{k}),\tilde{z}_{k}-z_{r}\rangle+\tfrac{\xi}{\sqrt{\lambda}}\cdot 2\beta_{k}^{\prime}+\tfrac{2\zeta}{1-\gamma}.$ (101)
Where we used
$\left\lVert{x}\right\rVert_{V_{k-1}^{-1}}=\sqrt{x^{\top}V_{k-1}^{-1}x}\leq\sqrt{\lambda_{\max}(V_{k-1}^{-1})x^{\top}x}\leq\tfrac{1}{\sqrt{\lambda}}\left\lVert{x}\right\rVert_{2}$ (102)
in the last step. [^8] This approximation of $\lambda_{\max}(V_{k-1}^{-1})\leq 1/\lambda$ is rather crude and could be improved by investigating how the eigenvalues of $V_{k}$ evolve over time.
We can treat the added term $2\beta_{k}^{\prime}\tfrac{\xi}{\sqrt{\lambda}}+\tfrac{2\zeta}{1-\gamma}$ separately throughout, bounding it by
$2\beta_{n}^{\prime}n\tfrac{\xi}{\sqrt{\lambda}}+\tfrac{n\zeta}{1-\gamma}$ (103)
after summing over the episodes. As a result, inserting our new $\beta_{n}^{\prime}$ in the regret bound, we obtain
$R_{n}$ $\leq 2\alpha\left(\sqrt{\lambda}S+\sigma\sqrt{d\log\left(\frac{d\lambda+nHL^{2}}{d\lambda}\right)+2\log n}+\underbrace{\frac{\sqrt{2}L\zeta}{\sqrt{\lambda}}\sqrt{ndH\log\left(\frac{d\lambda+nHL^{2}}{d\lambda}\right)}}_{\text{new}}\right)$
$\times\left(\sqrt{2nd\log\left(\frac{d\lambda+n\frac{L^{2}(1-\gamma^{H})^{2}}{(1-\gamma)^{2}}}{d\lambda}\right)}+\underbrace{\tfrac{2\xi n}{\sqrt{\lambda}}}_{\text{new}}\right)$
$+\underbrace{\tfrac{n\zeta}{1-\gamma}}_{\text{new}}$ (104)
Because of the new term appearing in $\beta^{\prime}$ and the final constant term, the regret bound is no longer sublinear, but instead
$R_{n}\leq\tilde{\mathcal{O}}(d\sqrt{n})+\zeta\tilde{\mathcal{O}}(nd)+\xi\tilde{\mathcal{O}}(n\sqrt{dn})$ (105)
which is also the final regret when incorporating the event $\mathcal{E}^{C}$. This bound indicates that, while the confidence set may shrink initially, the two mismatches may introduce an irreducible uncertainty in the size of the respective missmatches that cannot be resolved. Crucially, the suboptimality incurred is directly connected to the degree of mismatch considered in assumptions (A1’) and (A2’).
A.6 Useful Properties
A.6.1 Determinants